# Agentic AI Software Engineers: Programming with Trust

### v3

@@ Abstract.
Large Language Models (LLMs) have shown surprising proficiency in generating code snippets, promising to automate large parts of software engineering via artificial intelligence (AI). We argue that successfully deploying AI software engineers requires a level of trust equal to, or even greater than, the trust established by human-driven software engineering practices. The recent trend toward LLM agents offers a path toward integrating the power of LLMs to create new code with the power of analysis tools to increase trust in the code. This opinion piece comments on how LLM agents present AI software engineering workflows of the future, and whether the focus of programming will shift from scale to trust.
@@ 1. A Key Barrier for AI Software Engineers
This brings us to the following question: What is the place of AI software engineers in future development workflows? If we can figure out how automatically generated and manually written software can co-exist, this may give us a pathway of greater deployment of AI in software engineering! Starting from early programs of just a few lines written in high-level languages in the 1950s, the size of programs has increased greatly to hundreds of millions of lines of code. For the past fifty years, there has been a steady interest towards programming in the large, where distributed software development by different teams is carried out across different modules of a software project. With the increased use of AI code generation, we believe that the emphasis will be not only on programming at scale, but increasingly on programming with trust.
@@ 2. Technical and Human Trust
| Aspect | Trust Factor |
| Technical | Correctness: Does the code produce the expected results? |
| Security: Is the code free of vulnerabilities and safe to deploy? | |
| Performance: Is the code efficient and scalable? | |
| Maintainability: Is the code easy to read, refactor, and extend? | |
| Compliance: Does the code adhere to language/framework best practices? | |
| Testing: Are there sufficient and meaningful test cases? | |
| Dependencies: Are dependencies safe, current, and widely trusted? | |
| Human | Explainability and Transparency: Can the AI justify the chosen solution? |
| Bias and Ethics: Does the code reflect unbiased, ethical decisions? | |
| Familiarity: Are the code style or algorithms familiar to the developer? | |
| Over-Reliance: Is there concern the user might use the code blindly? | |
| Experience Mismatch: Does the code align with the developer’s expertise level? | |
| Collaboration: Does the AI react well to developer feedback? | |
| Team Practice: Does the AI align with the team’s development workflow, conventions, and review culture? |
Table 1 illustrates how developers perceive trust in AI-generated code, both from a technical and a human angle. Technical trust in AI-generated code stems from measurable attributes that determine code quality and reliability. Developers need assurance that the code is correct, secure, and performs efficiently under expected conditions. Maintainability and standards compliance are crucial for long-term usability, while thorough testing validates that the code functions as intended. Additionally, trust improves when dependencies are well-managed and sourced from reputable libraries.
Human trust involves psychological and social dynamics that influence how developers perceive and accept AI-generated solutions. A key factor is explainability and transparency: When an AI system can clarify its reasoning and design choices, developers are more likely to trust and adopt its outputs. Developers are also more likely to trust code that aligns with ethical values, follows familiar patterns and team practices, and matches the developers level of expertise. An important risk that could reduce trust is over-reliance on AI-provided code without properly understanding it. Furthermore, trust is strengthened when AI systems collaborate effectively and adapt to feedback provided by developers.
@@ 3. Software Engineering Agents
Recently, several software engineering agents have been proposed, starting with the announcement of the Devin AI software engineer from Cognition Labs (Wu, 2024). Devin can solve natural language tasks (called issues), such as bug fixes and feature additions. It combines a back-end LLM with access to standard developer tools, such as a shell, a code editor, and a web browser. The agent employs such tools autonomously to let the AI software engineer mimic human practices. In parallel with the announcement of Devin, several research groups proposed their own LLM agents for software engineering, including RepairAgent (Bouzenia et al., 2025), AutoCodeRover (Zhang et al., 2024) and SWE-agent (Yang et al., 2024). RepairAgent (Bouzenia et al., 2025) fixes bugs based on test cases, and guides the agent by defining a finite-state machine that outlines the typical steps followed by a developer. Constraining the agent’s autonomy in this way can increase trust by aligning the agent with the typical workflow followed by a human developer. AutoCodeRover, a spinoff acquired by SonarSource, can solve natural language issues requiring bug fixing or feature addition. It establishes technical trust, e.g., by using program analysis on abstract syntax trees, and human trust, e.g., by extracting the intent of the software, which can be used to provide explanations of the suggested code edits. SWE-agent (Yang et al., 2024) follows a philosophy similar to Devin, by making file navigation tools and interfaces available to an AI software engineer. It works on a text-based program representation and does not employ any program analysis.
@@ 4. Establishing Trust
What makes us trust human-written code, but not necessarily the code generated by an LLM? Part of the reason is the perceived capability of “passing the blame”. If a human developer is involved, there is the promise of getting feedback from the developer as needed. Of course, this does not always hold, e.g., if the developer eventually leaves an organization. Nevertheless, accepting a code commit from a developer often partially depends on the reputation of the developer within the organization. For an AI software engineer in the form of an LLM agent to earn a reputation similar to a senior human colleague, it has to integrate established quality assurance techniques and collaborate effectively with human developers. We outline several ideas toward this goal.
@@
A key factor for increasing human trust into AI software engineers will be to enforce effective AI-human collaboration patterns. As experienced first-hand by the AutoCodeRover team (Zhang et al., 2024) via real-life anecdotes communicated by clients, developer hesitation in accepting AI-generated code also comes from the volume of code that can be quickly generated by AI tools, overwhelming human developers. Supporting AI generated code with confidence scores can help reduce developer hesitation. Moreover, personalizing LLM agents to the experiences and expertise of specific users can help in increasing familiarity with AI-suggested code. The latest release of the Codex agent (OpenAI, 2025) in May 2025 (the Codex agent can run synchronously in the terminal watching a human, while also running long-running tasks in the cloud asynchronously) sharpens these questions – can a LLM coding agent be trusted as an intern or a junior team member?
Realistic Expectations and Critical Assessment
Clear documentation of an agent’s capabilities and known limitations can support informed decision making and can set realistic expectations. Any social biases being implicitly introduced by agents will need to be assessed. Integrating such assessments into the day-to-day development workflows enables teams to continuously monitor and improve the trustworthiness.
@@ 5. Outlook
As AI software engineers take on more core development tasks, their success will hinge not just on technical capability but on earning developer trust. Rather than relying on many separate agents for specialized software engineering tasks, it could be worthwhile to create a unified software engineering agent that combines coding, testing, debugging, etc. into a coherent, explainable workflow. To become a trusted collaborator, such an agent must offer transparency, adapt to feedback, and integrate safeguards that ensure quality and security. Ultimately, programming with AI will mean not full automation, but effective delegation – where human oversight and AI support work hand in hand.
Acknowledgements.
The authors acknowledge discussions with attendees at Dagstuhl Seminar 24431 on Automated Programming and Program Repair (Oct 2024), as well as at Shonan Meeting 217 on Trusted Automatic Programming (Jan 2025). Lingming Zhang shared valuable pointers.References
- (1)
- Bouzenia et al. (2025) Islem Bouzenia, Premkumar Devanbu, and Michael Pradel. 2025. Repairagent: An autonomous, LLM-based agent for program repair. In 47th International Conference on Software Engineering (ICSE).
- Chakraborty et al. (2025) Saikat Chakraborty, Gabriel Ebner, Siddharth Bhat, Sarah Fakhoury, Sakina Fatima, Shuvendu Lahiri, and Nikhil Swamy. 2025. Towards Neural Synthesis for SMT-assisted Proof-Oriented Programming . In 2025 IEEE/ACM 47th International Conference on Software Engineering (ICSE).
- Kohler (2025) Lindsay Kohler. 2025. Fear Isn’t Slowing Down AI Adoption At Work. A Lack Of Vision Is. Forbes Newsletter (Jan 2025). https://www.forbes.com/sites/lindsaykohler/2025/01/27/fear-isnt-slowing-down-ai-adoption-at-work-a-lack-of-vision-is/.
- Miserendino et al. (2025) Samuel Miserendino, Michele Wang, Tejal Patwardhan, and Johannes Heidecke. 2025. SWE-Lancer: Can Frontier LLMs Earn $1 Million from Real-World Freelance Software Engineering? arXiv:2502.12115 (Feb 2025).
- OpenAI (2025) OpenAI. 2025. Codex and Codex CLI. https://openai.com/codex/.
- Ruan et al. (2025) Haifeng Ruan, Yuntong Zhang, and Abhik Roychoudhury. 2025. SpecRover: Code Intent Extraction via LLMs. In 47th International Conference on Software Engineering (ICSE).
- Ryan et al. (2024) Gabriel Ryan, Siddhartha Jain, Mingyue Shang, Shiqi Wang, Xiaofei Ma, Murali Krishna Ramanathan, and Baishakhi Ray. 2024. Code-aware prompting: A study of coverage-guided test generation in regression setting using LLM. Proceedings of the ACM on Software Engineering 1, FSE (2024), 951–971.
- Wu (2024) Scott Wu. 2024. Introducing Devin, the first AI software engineer. Cognition Labs blogpost (Mar 2024). https://www.cognition.ai/blog/introducing-devin.
- Yang et al. (2024) John Yang, Carlos E. Jimenez, Alexander Wettig, Kilian Lieret, Shunyu Yao, Karthik Narasimhan, and Ofir Press. 2024. SWE-agent: Agent-computer Interfaces Enable Automated Software Engineering. Thirty-Eighth Annual Conference on Neural Information Processing Systems (NeurIPS) (2024).
- Zhang et al. (2024) Yuntong Zhang, Haifeng Ruan, Zhiyu Fan, and Abhik Roychoudhury. 2024. AutoCodeRover: Autonomous Program Improvement. In ACM International Symposium on Software Testing and Analysis (ISSTA).

### v4

@@ Abstract.
Large Language Models (LLMs) have shown surprising proficiency in generating code snippets, promising to automate large parts of software engineering (SE) via artificial intelligence (AI). We argue that successfully deploying AI software engineers requires a level of trust equal to, or even greater than, the trust established by human-driven software engineering practices. The recent trend toward LLM agents offers a path toward integrating the power of LLMs to create new code with the power of analysis tools to increase trust in the code. This opinion piece comments on how LLM agents present AI software engineering workflows of the future, and whether the focus of programming will shift from scale to trust.
@@ 1. A Key Barrier for AI Software Engineers
This brings us to the following question: What is the place of AI software engineers in future development workflows? If we can figure out how automatically generated and manually written software can co-exist, this may give us a pathway of greater deployment of AI in software engineering! Starting from early programs of just a few lines written in high-level languages in the 1960s and 70s, the size of programs has increased greatly to hundreds of millions of lines of code. For the past fifty years, there has been a steady interest towards programming in the large. With the increased use of AI code generation, we believe that the emphasis will be not only on programming at scale, but increasingly on programming with trust.
@@ 2. Technical and Human Trust
| Aspect | Trust Factor (What) | Trust Factor (How) |
| Technical | Correctness: Code produce the expected results | Reviewing & Testing: Test the AI-generated code with sufficient and meaningful test cases |
| Security: Code is vulnerability-free and safe to deploy | Dependency analysis: Check whether the dependencies are safe, up-to-date, and trusted | |
| Performance: Code is efficient and scalable | Profiling: Measure performance during execution | |
| Maintainability: Code is easy to read, refactor, and extend | Metrics: Measure code complexity and readability | |
| Compliance: Code adheres to language/framework best practices | Static analysis: Check for rule violations, e.g., with linters | |
| Human | Explainability and Transparency: AI justifies the chosen solution | Familiarity: Check whether the code is explained in ways suitable for the developer |
| Bias and Ethics: Code reflects unbiased, ethical decisions | Over-reliance: Monitor and estimate if the developer is using AI coding assistance blindly | |
| Team Practice: AI aligns with the team’s development workflow, experience, and review culture | Experience mismatch: Check whether the AI generated code align with the developer’s expertise level | |
| Collaboration: Coding agent reacts well to developer feedback |
Table 1 illustrates how developers perceive trust in AI-generated code, both from a technical and a human angle. Technical trust in AI-generated code stems from measurable attributes that determine code quality and reliability. Developers need assurance that the code is correct, secure, and performs efficiently under expected conditions. Maintainability and standards compliance are crucial for long-term usability, while thorough testing validates that the code functions as intended. Additionally, trust improves when dependencies are well-managed and sourced from reputable libraries. Human trust involves psychological and social dynamics that influence how developers perceive and accept AI-generated solutions. A key factor is explainability and transparency: when an AI system can clarify its reasoning and design choices, developers are more likely to trust and adopt its outputs. Developers are also more likely to trust code that aligns with ethical values, follows familiar patterns and team practices, and matches the developers level of expertise. Trust is strengthened when AI systems collaborate effectively and adapt to feedback provided by developers.
Even though we distinguish technical and human trust, we do not attempt to relate to interpersonal trust models from sociology and psychology. In fact, we feel that agent-developer interactions in software will be a new mode of interaction not seen before, and new trust models may develop, e.g., building on prior efforts toward modeling trust in human-AI interactions (Mehrotra et al., 2024).
@@ 3. Software Engineering Agents
Recently, several software engineering agents have been proposed, starting with the announcement of the Devin AI software engineer from Cognition Labs (Wu, 2024). Devin can solve natural language tasks (called issues), such as bug fixes and feature additions. It combines a back-end LLM with access to standard developer tools, such as a shell, a code editor, and a web browser. The agent employs such tools autonomously to let the AI software engineer mimic human practices. In parallel with the announcement of Devin, several research groups proposed their own LLM agents for software engineering, including RepairAgent (Bouzenia et al., 2025), AutoCodeRover (Zhang et al., 2024) and SWE-agent (Yang et al., 2024). RepairAgent (Bouzenia et al., 2025) fixes bugs exposed by failing test cases, and guides the agent by defining a finite-state machine that outlines the typical steps followed by a developer. RepairAgent can only work with tests and cannot process natural language issues. AutoCodeRover, a spinoff acquired by SonarSource, can solve natural language issues requiring bug fixing or feature addition. It has been integrated into the widely used SonarQube static analyzer and has already been made available to enterprise customers. It establishes technical trust, e.g., by using program analysis on abstract syntax trees, and human trust, e.g., by extracting the intent of the software, which can be used to provide explanations of the suggested code edits. However, it makes less use of file navigation and bash tools in its implementation. SWE-agent (Yang et al., 2024) follows a philosophy similar to Devin, by making file navigation tools and interfaces available to an AI software engineer. It does not employ any program analysis and hence cannot do intent extraction.
@@ 4. Establishing Trust
What makes us trust human-written code, but not necessarily the code generated by an LLM? Part of the reason is the perceived capability of “passing the blame”. If a human developer is involved, there is the promise of getting feedback from the developer as needed. Of course, this does not always hold, e.g., if the developer eventually leaves an organization. Nevertheless, accepting a code commit from a developer partially depends on the reputation of the developer within the organization. For an AI software engineer in the form of an LLM agent to earn a reputation similar to a senior human colleague, it has to integrate established quality assurance techniques and collaborate effectively with human developers. We outline several ideas toward this goal.
@@
A key factor for increasing human trust into AI software engineers will be to enforce effective AI-human collaboration patterns. As experienced first-hand by the AutoCodeRover team (Zhang et al., 2024) via real-life anecdotes communicated by clients, developer hesitation in accepting AI-generated code also comes from the volume of code that can be quickly generated by AI tools, overwhelming human developers. Providing AI generated code with confidence scores can reduce developer hesitation. Validating (or refuting) such anecdotal experiences via studies on how to reduce developer hesitation remains a direction of future research. Latest release of OpenAI’s Codex agent in May 2025 (the Codex agent can run synchronously in the terminal watching a human, while also running long-running tasks in the cloud asynchronously) sharpens these questions – can a LLM coding agent be trusted? To operationalize socio-technical integration for AI agent trust, organizations should implement review parity—requiring AI-generated code changes to undergo identical peer review processes as human contributions, e.g., by enforcing the same two-reviewer gates and quality thresholds. Additionally, deploying specialized code review agents backed by trusted static and dynamic analysis tools may assist human reviewers. In the development workflow, organizations should incorporate transparent scaffolding by systematically tagging AI-generated code, embedding authorship provenance and confidence metrics directly within pull request templates and review interfaces. In this way, trust measures can be integrated inside software workflows.
@@ 5. Outlook
As AI software engineers take on more core development tasks, their success will hinge not just on technical capability but on earning developer trust. Rather than relying on many separate agents for specialized software engineering tasks, it could be worthwhile to create a unified software engineering agent that combines coding, testing, debugging, etc. into a coherent, explainable workflow. To become a trusted collaborator, such an agent must offer transparency, adapt to feedback, and integrate safeguards that ensure quality and security. Programming with AI will mean not full automation, but effective delegation, where human and AI work hand in hand.